OpenStudy (anonymous):

What is the equation for a line that contains the point (-4,-1) (3,-3)

OpenStudy (anonymous):

here's how to solve all of these types of questions. Calculate the slope of the line by doing change in y over change in x, or (y2 - y1)/(x2 - x1). Then use "y = mx + b". Plug in the slope (m) and one of your points for x and y. Solve for b. Then you have m and b, and the equation of the line is y = mx + b.

OpenStudy (anonymous):

SO WOULD IT BE -2/7 - 2x

OpenStudy (anonymous):

m is (-3 - -1)/(3 - -4) = -2/7 now y = mx + b, and plug in the first point: -1 = (-2/7)(-4) + b -1 = 8/7 + b -15/7 = b Then the equation: y = -2/7 x - 15/7